OpenStudy (anonymous):

dividing monimials 7c^7/-7c^2

OpenStudy (mathstudent55):

Divide the number parts: 7/(-7) To divide the variables, use the rule of dividing powers: a^m/a^n = a^(m - n) For example: 28a^5/-4a^2 = -7a^3

OpenStudy (anonymous):

7 and 7 get cancelled, but the negative sign remains... then since c is common, u subtract the powers...7-2 = c^5..

OpenStudy (anonymous):

-c^5

OpenStudy (anonymous):

-c^5 is the right answer